# DynGMA: a robust approach for learning stochastic differential equations from data

### v1

@@ 1 Introduction
Dynamical systems with random perturbations can be described by Stochastic Differential Equations (SDE) driven by Brownian noise. This modeling approach serves as a scientific modeling tool applied across a diverse array of disciplines, such as biology, physics, and chemistry. In practical scenarios, the underlying dynamics, including the drift and diffusion functions, are often unknown or only partially understood, while a finite set of discrete observational data is available. For example, various physical experimental systems generate data pertaining to motion tracking chen2023constructing ; chen2022automated ; schmidt2009distilling , microscopic fine-gained physics simulations yield macroscopic coarse-grained statistics chen2023constructing ; dietrich2023learning and the concentrations of various individuals in an epidemic process can be obtained through Monte Carlo simulations dietrich2023learning ; makeev2002coarse . In these cases, researchers have to discover interpretable SDE models from data and then infer information such as future behavior, invariant distribution, and energy landscape of the system.
@@
The discovery of dynamical systems, including systems with randomness, from observed data is a well-established task bongard2007automated ; brunton2017chaos ; rudy2017data ; schmidt2009distilling . SDEs are characterized by their transition densities, which are typically challenging to express in a closed form due to their involvement in solving the Fokker-Planck-Kolmogorov (FPK) equation. As the transition density is essential for likelihood-based parametric estimation of SDEs sampled at discrete times, various approximations have been proposed, e.g., sampling-based approaches brandt2002simulated ; pedersen1995new , methods based on approximate solutions to the FPK equation jensen2002transition , higher-order Ito-Taylor expansion method kessler1997estimation , and Hermite polynomials expansion method ait2002maximum . Details of the methods on this subject can be found in monographs iacus2008simulation ; prakasa1999statistical . Recently, thanks to the exponential growth of available data and computing resources, machine learning has emerged as a flexible and data-driven approach for learning fully unknown stochastic dynamics chen2021solving ; chen2023constructing ; chen2023learning ; dietrich2023learning ; gu2023stationary ; look2022deterministic ; solin2021scalable ; xu2023modeling . In addition, SDEs are also employed as a black box model in a diverse range of machine learning tasks li2020scalable ; look2019differential ; song2021score ; tzen2019neural .
@@
It has been shown that sampling-based approaches converge to the true transition density as the number of samples increases brandt2002simulated ; pedersen1995new . However, such approaches are typically used for small-scale parameter scenarios where the dynamics is not fully unknown, and require drawing a large number of sample trajectories to accurately represent the underlying distribution as discussed and numerically validated in look2022deterministic . These issues are addressed by developing deterministic Gaussian approximations to the transition density look2022deterministic ; solin2021scalable . Specifically, Look et al. look2022deterministic propose bidimensional moment matching (BMM) to approximate the transition density by a Gaussian density, relying on the layered structure of neural networks. However, the dependence on the layered structure limits its flexibility and makes it difficult to encode partially known physical terms. In addition, Solin et al. solin2021scalable propose FPK moment matching (FPK-MM) by Gaussian assumed density approximations of the FPK equation following classical SDE techniques kushner1967approximations ; sarkka2013gaussian . However, both approaches treat SDEs as a black box and have not thoroughly investigated its performance in learning interpretable governing functions of SDE models. Moreover, the Gaussian approximation and the exact density function evidently do not coincide, and the error quantification of existing Gaussian-based approaches remains open.
@@
The most straightforward and efficient Gaussian approximation can be obtained by employing the one-step Euler scheme to the Gaussian approximation of the FPK equation. Alternatively, it can also be obtained through the one-step Euler-Maruyama scheme dietrich2023learning . We refer to this approach as the Euler-Maruyama approach in this paper. From this point of view of stochastic numerical integrators, approximations of the transition density beyond Gaussian are derived for other one-step numerical schemes, including Milstein and Runge-Kutta type integrators dietrich2023learning . The error of such approximations can thus be deduced from error analysis of the underlying numerical schemes, and numerical results demonstrate the effectiveness of learning SDEs using the proposed approximations chen2023constructing ; dietrich2023learning ; lin2023computing . In addition, Gu et al. gu2023stationary derive the loss function directly rather than using likelihood loss, also based on the Euler-Maruyama scheme. However, methods in this direction may have the limitation of requiring data with sufficiently high time resolution since they are all based on one-step numerical schemes, which will be demonstrated in our numerical experiments later.
@@
Table 1: Comparison between Sampling-based method brandt2002simulated ; pedersen1995new , BMM look2022deterministic , FPK-MM solin2021scalable , Integrators-based method dietrich2023learning and DynGMA. The symbol $\checkmark$ indicates that the method maintains reliability and good performance under these scenarios, or possesses these properties.
@@
Sampling-based brandt2002simulated ; pedersen1995new partially
BMM look2022deterministic
FPK-MM kushner1967approximations ; sarkka2013gaussian ; solin2021scalable partially
Integrators-based dietrich2023learning
@@
In this paper, we employ neural networks to parameterize the drift and diffusion functions of the fully unknown SDE, and use likelihood-based loss to train the network. By revisiting the random perturbation theory of dynamical systems blagoveshchenskii1962diffusion ; blagoveshchenskii1961certain , we propose a novel Gaussian density approximation, and its extension, named as dynamical Gaussian mixture approximation (DynGMA), to approximate the transition densities of the parameterized SDE. Specifically, we discretize the time step into several sub-intervals, and employ the proposed Gaussian approximation in each sub-interval. We quantify the error of the proposed Gaussian approximation, indicating the advantages of our method compared to the Gaussian approximation obtained via the Euler-Maruyama scheme. With DynGMA, we can obtain the automatically-differentiable approximation of the fully unknown drift and diffusion functions of the SDE, and we can then compute the invariant distribution of unknown stochastic dynamics from trajectory data that does not necessarily follow the equilibrium distribution. Moreover, our method operates independently of specific structures imposed on stochastic dynamics and thus naturally accommodates the encoding of partially known physical terms, such as generalized Onsager principle chen2023constructing ; yu2021onsagernet , symplectic structure greydanus2019hamiltonian , and GENERIC formalism zhang2022gfinns . In addition to possessing these advantages, the notable benefit of our approach is its ability to effectively learn SDEs even when using trajectory data with low time resolution. This property enables DynGMA to incorporate multi-step training williams1989learning and be extended to a recurrent version without any modification, allowing it to effectively handle data with measurement noise. A detailed comparison between our method and the aforementioned approaches is presented in the Table 1.
@@ 2 Learning SDEs from data and its applications
where $\Delta t_{m}=t_{m}-t_{m-1}$. We typically require an approximation of the transition density, denoted as $p_{h}(\Delta t,x|0,x_{0})$. For example, according to the Euler-Maruyama scheme, we can obtain the most straightforward and efficient Gaussian approximation dietrich2023learning :
@@
In addition, if an approximation of the transition density is also valid for multiple steps, we instead introduce the multi-step training, a well-known technique in time series analysis williams1989learning , by maximizing the likelihood objective:
@@
As the transition density is essential for the MLE and existing methods mostly rely on Gaussian approximations dietrich2023learning ; look2022deterministic ; solin2021scalable or one-step numerical schemes dietrich2023learning , we will quantify the errors associated with Gaussian approximations based on the random perturbation theory of dynamical systems blagoveshchenskii1962diffusion ; blagoveshchenskii1961certain , and develop novel approximations that allow moderately large time steps in the next section.
@@ 3.1 Dynamical Gaussian mixture approximation
Existing methods are usually based on one-step stochastic numerical schemes and thus necessitate data with sufficiently high time resolution. To make multiple time steps valid, thereby obtaining a more accurate approximation for large time intervals, we use the Gaussian mixture model to approximate the transition density $p(\Delta t,x(\Delta t)|0,x(0);\theta)$ of SDE (3) in this section. We refer to this approximation as dynamical Gaussian mixture approximation (DynGMA).
@@
Subsequently, using Chapman-Kolmogorov Equation pavliotis2016stochastic and the Markov property of $\{x(kh)\}_{k=0}^{K}$, we are able to write the $k$-step transition density:
@@
$\int\mathcal{N}(x_{k+1}|\mu(x_{k}),\Sigma(x_{k}))\mathcal{N}(x_{k}|\mu^{i}_{k},\Sigma^{i}_{k})\mathop{}\!\mathrm{d}x_{k}\approx\sum_{j=0}^{J}v^{j}\mathcal{N}(x_{k+1}|\mu(\mu^{i}_{k}+\sqrt{\Sigma^{i}_{k}}{\xi}_{j}),\Sigma(\mu^{i}_{k}+\sqrt{\Sigma^{i}_{k}}{\xi}_{j})),$
where $v^{j}$ and ${\xi}_{j}$ are predetermined coefficients. Subsequently, we obtain the Gaussian mixture approximation of $p((k+1)h,x_{k+1}|0,x_{0})$, yielding $p_{h}((k+1)h,x_{k+1}|0,x_{0})$:
@@ 3.2 Gaussian density approximation via the asymptotic expansion of SDEs
As mentioned in the introduction, there are various Gaussian density approximations to the transition density. Since the error estimation of the existing Gaussian approximations is either lacking or based on one-step numerical schemes, in this section, we propose a novel Gaussian density approximation inspired directly by the random perturbation theory of dynamical systems blagoveshchenskii1962diffusion ; blagoveshchenskii1961certain . The proposed Gaussian density approximation is derived via a continuous ordinary differential equation. We have the flexibility to choose discretization methods (including adjusting the number of steps) in order to attain higher accuracy. Based on the random perturbation theory, we will also be able to quantify the error of the proposed Gaussian approximations and demonstrate the advantages of our method compared to the Euler-Maruyama approach (5). The proofs are postponed to C.
Next, we present the theorem of Blagoveshchenskii blagoveshchenskii1962diffusion , which we further extend to learning SDEs using neural networks.
@@
where $J_{f}=\frac{\partial f}{\partial x}$ denotes the Jacobian of $f$, and E[]normal-$\mathop{}\!\mathrm{E}[\cdot]$ denotes the expectation over the Brownian motion.
@@
where $J_{f_{\theta}}=\frac{\partial f_{\theta}}{\partial x}$ represents the Jacobian of $f_{\theta}$, and E[]normal-$\mathop{}\!\mathrm{E}[\cdot]$ denotes the expectation over the Brownian motion.
@@
Input time step $\Delta t\geq 0$, initial value $x_{0}\in\mathbb{R}^{D}$ and neural networks $f_{\theta}$ and $\sigma_{\theta}$.
@@
Input time step $\Delta t\geq 0$, initial value $x_{0}\in\mathbb{R}^{D}$ and neural networks $f_{\theta}$ and $\sigma_{\theta}$.
@@
$I_{1}=1$, $\omega_{1}^{0}=1$.
$\mu_{1}^{0}=\mu(x_{0})$, $\sqrt{\Sigma_{1}^{0}}=\sqrt{\Sigma(x_{0})}$ according to (15).
for $11\leq k\leq K-$ 1 do
@@
The proposed DynGMA is closely related to the Euler-Maruyama approach dietrich2023learning . In the case when there is solely one Gaussian mixture step (i.e., $K=1$), and the selected discretization scheme for (13) is the one-step Euler method, DynGMA transforms into a Gaussian approximation (5) derived through the Euler-Maruyama approach. According to the random perturbation theory of dynamical systems, the estimates for (5) can be written as follows:
@@ 3.3 Parameterization and loss function
the proposed approximation $p_{h}(\Delta t,x_{K}|0,x_{0};\theta)$ has a smooth representation, and thus has forward and backward implementations in established automatic differentiation packages. We can choose to optimize the loss using stochastic gradient-based optimization methods such as Adam optimization kingma2014adam . Alternatively, we propose multi-step training with DynGMA by optimizing the likelihood loss:
@@ 4 Numerical experiments
The baseline method used for comparison in this section is the Gaussian approximation (5) obtained via the Euler-Maruyama dietrich2023learning , of which effectiveness has been illustrated in various applications such as learning effective SDEs from microscopic simulations dietrich2023learning , computing invariant distributions from data lin2023computing , and learning macroscopic thermodynamic description of stretching dynamics of polymer chains chen2023constructing . We employ a similar parameterization for the baseline method, and its loss function is obtained by substituting (5) into (17), as follows:
@@
We refer to this approach as the Euler-Maruyama approach. Another method for comparison is the Gaussian approximation obtained via the FPK equation and the cubature approximation kushner1967approximations ; sarkka2013gaussian ; solin2021scalable . The corresponding loss function is of the form
@@
We refer to this approach as the Gaussian Cubature approach. Unless otherwise stated, the coefficients $v^{j}$ and ${\xi}_{j}$ in the above formula and in the DynGMA are both selected by the unscented transform arasaratnam2009cubature ; mcnamee1967construction :
@@
where $e_{j}$ is the unit coordinate vector. If not otherwise specified, we use the Euler-Maruyama scheme with very fine step size [^1] Specifically, $\Delta t/10$ in Section 4.2, $\Delta t/20$ in Section 4.3, $5\times 10^{-4}$ in Section 4.4, and $10^{-4}$ in Section 4.5. to simulate the underlying stochastic dynamics for generating data and to simulate the learned systems for numerical verification. Moreover, we compute the average relative error to provide a quantitative evaluation of the performance

### v2

@@ 1 Introduction
Dynamical systems with random perturbations can be described by Stochastic Differential Equations (SDE) driven by Brownian noise. This modeling approach serves as a scientific modeling tool applied across a diverse array of disciplines, such as biology, physics, and chemistry. In practical scenarios, the underlying dynamics, including the drift and diffusion functions, are often unknown or only partially understood, while a finite set of discrete observational data is available. For example, various physical experimental systems generate data pertaining to motion tracking [11, 10, 41], microscopic fine-gained physics simulations yield macroscopic coarse-grained statistics [11, 14] and the concentrations of various individuals in an epidemic process can be obtained through Monte Carlo simulations [14, 33]. In these cases, researchers have to discover interpretable SDE models from data and then infer information such as future behavior, invariant distribution, and energy landscape of the system.
@@
The discovery of dynamical systems, including systems with randomness, from observed data is a well-established task [6, 8, 38, 41]. SDEs are characterized by their transition densities, which are typically challenging to express in a closed form due to their involvement in solving the Fokker-Planck-Kolmogorov (FPK) equation. As the transition density is essential for likelihood-based parametric estimation of SDEs sampled at discrete times, various approximations have been proposed, e.g., sampling-based approaches [7, 36], methods based on approximate solutions to the FPK equation [21], higher-order Ito-Taylor expansion method [22], and Hermite polynomials expansion method [1]. Details of the methods on this subject can be found in monographs [20, 37]. Recently, thanks to the exponential growth of available data and computing resources, machine learning has emerged as a flexible and data-driven approach for learning fully unknown stochastic dynamics [12, 11, 13, 14, 16, 31, 42, 47]. In addition, SDEs are also employed as a black box model in a diverse range of machine learning tasks [28, 30, 43, 45].
@@
It has been shown that sampling-based approaches converge to the true transition density as the number of samples increases [7, 36]. However, such approaches are typically used for small-scale parameter scenarios where the dynamics is not fully unknown, and require drawing a large number of sample trajectories to accurately represent the underlying distribution as discussed and numerically validated in [31]. These issues are addressed by developing deterministic Gaussian approximations to the transition density [31, 42]. Specifically, Look et al. [31] propose bidimensional moment matching (BMM) to approximate the transition density by a Gaussian density, relying on the layered structure of neural networks. However, the dependence on the layered structure limits its flexibility and makes it difficult to encode partially known physical terms. In addition, Solin et al. [42] propose FPK moment matching (FPK-MM) by Gaussian assumed density approximations of the FPK equation following classical SDE techniques [25, 39]. However, both approaches treat SDEs as a black box and have not thoroughly investigated its performance in learning interpretable governing functions of SDE models. Moreover, the Gaussian approximation and the exact density function evidently do not coincide, and the error quantification of existing Gaussian-based approaches remains open.
@@
The most straightforward and efficient Gaussian approximation can be obtained by employing the one-step Euler scheme to the Gaussian approximation of the FPK equation. Alternatively, it can also be obtained through the one-step Euler-Maruyama scheme [14]. We refer to this approach as the Euler-Maruyama approach in this paper. From this point of view of stochastic numerical integrators, approximations of the transition density beyond Gaussian are derived for other one-step numerical schemes, including Milstein and Runge-Kutta type integrators [14]. The error of such approximations can thus be deduced from error analysis of the underlying numerical schemes, and numerical results demonstrate the effectiveness of learning SDEs using the proposed approximations [11, 14, 29]. In addition, Gu et al. [16] derive the loss function directly rather than using likelihood loss, also based on the Euler-Maruyama scheme. However, methods in this direction may have the limitation of requiring data with sufficiently high time resolution since they are all based on one-step numerical schemes, which will be demonstrated in our numerical experiments later.
@@
Table 1: Comparison between Sampling-based method [7, 36], BMM [31], FPK-MM [42], Integrators-based method [14] and DynGMA. The symbol $\checkmark$ indicates that the method maintains reliability and good performance under these scenarios, or possesses these properties.
@@
Sampling-based [7, 36] partially
BMM [31]
FPK-MM [25, 39, 42] partially
Integrators-based [14]
@@
In this paper, we employ neural networks to parameterize the drift and diffusion functions of the fully unknown SDE, and use likelihood-based loss to train the network. By revisiting the random perturbation theory of dynamical systems [4, 5], we propose a novel Gaussian density approximation, and its extension, named as dynamical Gaussian mixture approximation (DynGMA), to approximate the transition densities of the parameterized SDE. Specifically, we discretize the time step into several sub-intervals, and employ the proposed Gaussian approximation in each sub-interval. We quantify the error of the proposed Gaussian approximation, indicating the advantages of our method compared to the Gaussian approximation obtained via the Euler-Maruyama scheme. With DynGMA, we can obtain the automatically-differentiable approximation of the fully unknown drift and diffusion functions of the SDE, and we can then compute the invariant distribution of unknown stochastic dynamics from trajectory data that does not necessarily follow the equilibrium distribution. Moreover, our method operates independently of specific structures imposed on stochastic dynamics and thus naturally accommodates the encoding of partially known physical terms, such as generalized Onsager principle [11, 48], symplectic structure [15], and GENERIC formalism [49]. In addition to possessing these advantages, the notable benefit of our approach is its ability to effectively learn SDEs even when using trajectory data with low time resolution. This property enables DynGMA to incorporate multi-step training [46] and be extended to a recurrent version without any modification, allowing it to effectively handle data with measurement noise. A detailed comparison between our method and the aforementioned approaches is presented in the Table 1.
@@ 2 Learning SDEs from data and its applications
where $\Delta t_{m}=t_{m}-t_{m-1}$. We typically require an approximation of the transition density, denoted as $p_{h}(\Delta t,x|0,x_{0})$. For example, according to the Euler-Maruyama scheme, we can obtain the most straightforward and efficient Gaussian approximation [14]:
@@
In addition, if an approximation of the transition density is also valid for multiple steps, we instead introduce the multi-step training, a well-known technique in time series analysis [46], by maximizing the likelihood objective:
@@
As the transition density is essential for the MLE and existing methods mostly rely on Gaussian approximations [14, 31, 42] or one-step numerical schemes [14], we will quantify the errors associated with Gaussian approximations based on the random perturbation theory of dynamical systems [4, 5], and develop novel approximations that allow moderately large time steps in the next section.
@@ 3.1 Dynamical Gaussian mixture approximation
Existing methods are usually based on one-step stochastic numerical schemes and thus necessitate data with sufficiently high time resolution. To make multiple time steps valid, thereby obtaining a more accurate approximation for large time intervals, we use the Gaussian mixture model to approximate the transition density $p(\Delta t,x(\Delta t)|0,x(0);\theta)$ of SDE (3) in this section. We refer to this approximation as dynamical Gaussian mixture approximation (DynGMA). An illustration of the proposed method is presented in Fig. 1.
Figure 1: Illustration of the proposed DynGMA with $3$ evaluation points. (Black box section) We discretize the time interval $[0,\Delta t]$ into $K$ sub-intervals with a step size of $h$, and then use Gaussian mixture model to approximate the $k$-step transition density (see Section 3.1 for details). (Red box section) At each step of size $h$, we solve (13) to determine each Gaussian density within the Gaussian mixture model (see Section 3.2 for details).
@@
Subsequently, using Chapman-Kolmogorov Equation [35] and the Markov property of $\{x(kh)\}_{k=0}^{K}$, we are able to write the $k$-step transition density:
@@
$\int\mathcal{N}(x_{k+1}|\mu(x_{k}),\Sigma(x_{k}))\mathcal{N}(x_{k}|\mu^{i}_{k},\Sigma^{i}_{k})\mathop{}\!\mathrm{d}x_{k}\approx\sum_{j=0}^{J}v^{j}\mathcal{N}(x_{k+1}|\mu(\mu^{i}_{k}+\sqrt{\Sigma^{i}_{k}}{\xi}_{j}),\Sigma(\mu^{i}_{k}+\sqrt{\Sigma^{i}_{k}}{\xi}_{j})).$
Here the coefficients ${\xi}_{j}$ and $v^{j}$ determine the evaluation points and corresponding weights, respectively. There are general settings independent of specific problems, which will be provided in Section 4. Subsequently, we obtain the Gaussian mixture approximation of $p((k+1)h,x_{k+1}|0,x_{0})$, yielding $p_{h}((k+1)h,x_{k+1}|0,x_{0})$:
@@ 3.2 Gaussian density approximation via the asymptotic expansion of SDEs
In this section, we derive the Gaussian approximation (9) employed in each sub-step. Due to the Markov property, it suffices to focus on the case where $k=0$. As mentioned in the introduction, there are various Gaussian density approximations to the transition density. Since the error estimation of the existing Gaussian approximations is either lacking or based on one-step numerical schemes, in this section, we propose a novel Gaussian density approximation inspired directly by the random perturbation theory of dynamical systems [4, 5]. The proposed Gaussian density approximation is derived via a continuous ordinary differential equation. We have the flexibility to choose discretization methods (including adjusting the number of steps) in order to attain higher accuracy. Based on the random perturbation theory, we will also be able to quantify the error of the proposed Gaussian approximations and demonstrate the advantages of our method compared to the Euler-Maruyama approach (5). The proofs are postponed to D.
Next, we present the theorem of Blagoveshchenskii [4], which we further extend to learning SDEs using neural networks.
@@
where $J_{f}=\frac{\partial f}{\partial x}$ denotes the Jacobian of $f$, and $\mathop{}\!\mathrm{E}[\cdot]$ denotes the expectation over the Brownian motion.
@@
where $J_{f_{\theta}}=\frac{\partial f_{\theta}}{\partial x}$ represents the Jacobian of $f_{\theta}$, and $\mathop{}\!\mathrm{E}[\cdot]$ denotes the expectation over the Brownian motion.
@@
Input time step $\Delta t\geq 0$, neural networks $f_{\theta}$ and $\sigma_{\theta}$.
Input initial mean and covariance $\mu_{0}^{0}$ and $\Sigma_{0}^{0}$ of $x_{0}$. If $x_{0}$ is a single point, then $\mu_{0}^{0}=x_{0}$ and $\Sigma_{0}^{0}=0$.
$I_{0}=1$, $w_{0}^{0}=1$.
@@
Input time step $\Delta t\geq 0$, neural networks $f_{\theta}$ and $\sigma_{\theta}$.
Input initial mean and covariance $\mu_{0}^{0}$ and $\Sigma_{0}^{0}$ of $x_{0}$. If $x_{0}$ is a single point, then $\mu_{0}^{0}=x_{0}$ and $\Sigma_{0}^{0}=0$.
$I_{0}=1$, $w_{0}^{0}=1$.
@@
for $1{0}\leq k\leq K-$ 1 do
@@
The proposed DynGMA is closely related to the Euler-Maruyama approach [14]. In the case when there is solely one Gaussian mixture step (i.e., $K=1$), and the selected discretization scheme for (13) is the one-step Euler method, DynGMA transforms into a Gaussian approximation (5) derived through the Euler-Maruyama approach. According to the random perturbation theory of dynamical systems, the estimates for (5) can be written as follows:
@@ 3.3 Parameterization and loss function
the proposed approximation $p_{h}(\Delta t,x_{K}|0,x_{0};\theta)$ has a smooth representation, and thus has forward and backward implementations in established automatic differentiation packages. We can choose to optimize the loss using stochastic gradient-based optimization methods such as Adam optimization [24]. Alternatively, we propose multi-step training with DynGMA by optimizing the likelihood loss:
@@ 4 Numerical experiments
The baseline method used for comparison in this section is the Gaussian approximation (5) obtained via the Euler-Maruyama [14], of which effectiveness has been illustrated in various applications such as learning effective SDEs from microscopic simulations [14], computing invariant distributions from data [29], and learning macroscopic thermodynamic description of stretching dynamics of polymer chains [11]. We employ a similar parameterization for the baseline method, and its loss function is obtained by substituting (5) into (17), as follows:
@@
We refer to this approach as the Euler-Maruyama approach. Another method for comparison is the Gaussian approximation obtained via the FPK equation and the cubature approximation [25, 39, 42]. The corresponding loss function is of the form
@@
We refer to this approach as the Gaussian Cubature approach. The computational complexity of both the Gaussian Cubature approach and DynGMA heavily relies on the selection of coefficients $\xi_{j}$ and $v^{j}$. A commonly chosen method is the symmetric cubature rule obtained through the unscented transform [2, 34, 40], which precisely encodes the given mean and covariance information in the selected set of points:
@@
where $e_{j}$ is the unit coordinate vector. This rule can be constructed up to any desired order [34]. Given our focus on dynamical systems with random perturbations, where the variance magnitude is typically small, we adopt the aforementioned parameter settings for both the Gaussian Cubature approach and DynGMA in this paper.
If not otherwise specified, we use the Euler-Maruyama scheme with very fine step size [^1] Specifically, $\Delta t/10$ in Section 4.2, $\Delta t/20$ in Section 4.3, $5\times 10^{-4}$ in Section 4.4, and $10^{-4}$ in Section 4.5. to simulate the underlying stochastic dynamics for generating data and to simulate the learned systems for numerical verification. Moreover, we compute the average relative error to provide a quantitative evaluation of the performance